# MDS-ED: Multimodal Decision Support

### v2

@@ Abstract
Background: Benchmarking medical decision support algorithms often struggles due to limited access to datasets, narrow prediction tasks, and restricted input modalities. These limitations affect their clinical relevance and performance in high-stakes areas like emergency care, complicating replication, validation, and improvement of benchmarks.
Methods: We introduce a dataset based on MIMIC-IV, benchmarking protocol, and initial results for evaluating multimodal decision support in the emergency department (ED). We use diverse data modalities from the first 1.5 hours of patient arrival, including demographics, biometrics, vital signs, lab values, and electrocardiogram waveforms. We analyze 1443 clinical labels across two contexts: predicting diagnoses with ICD-10 codes and forecasting patient deterioration.
Results: Our multimodal diagnostic model achieves an AUROC score over 0.8 in a statistically significant manner for 357 out of 1428 conditions, including cardiac issues like myocardial infarction and non-cardiac conditions such as renal disease and diabetes. The deterioration model scores above 0.8 in a statistically significant manner for 13 out of 15 targets, including critical events like cardiac arrest and mechanical ventilation, ICU admission as well as short- and long-term mortality. Incorporating raw waveform data significantly improves model performance, which represents one of the first robust demonstrations of this effect.
Conclusions: This study highlights the uniqueness of our dataset, which encompasses a wide range of clinical tasks and utilizes a comprehensive set of features collected early during the emergency after arriving at the ED. The strong performance, as evidenced by high AUROC scores across diagnostic and deterioration targets, underscores the potential of our approach to revolutionize decision-making in acute and emergency medicine.
@@ 1 Introduction
Artificial intelligence (AI) is an emerging field that has significantly enhanced healthcare and medicine [1] in areas such as precision medicine [2] and drug discovery [3]. The implementation of AI models is particularly relevant in acute and emergency medicine, where clinicians must address critical conditions within short time frames to make optimal clinical decisions. AI models can improve acute care by enabling early diagnosis [4], proposing tailored diagnostic workups [5], predicting admissions (including ward or intensive care unit (ICU)) [6], estimating survival rates [7], and providing faster and low-cost diagnoses [8]. There has been exponential growth in the number of scientific publications on new AI models. However, many of these studies have limitations, such as predicting only a limited set of specific conditions [9, 4], offering short time horizons for survival predictions, and requiring costly diagnostic tests [10].
@@
Healthcare datasets are essential for advancing medical research and innovation. However, much of this valuable data remains undisclosed due to commercial interests and privacy concerns. Despite these hurdles, there has been a recent surge in publicly accessible datasets aimed at facilitating the development and validation of machine learning models to tackle complex healthcare challenges [11, 12, 13, 14, 15]. Nonetheless, these datasets still have limitations, including size constraints [12], lack of open-source availability [16, 17], and narrow scopes in terms of prediction tasks [12, 13, 17, 14, 16, 15]. Benchmarking is crucial for evaluating the performance and reliability of novel AI-models, ensuring their applicability and robustness in real-world clinical settings. In this study, we investigate multimodal decision support for emergency department (MDS-ED) using an open-source biomedical multimodal dataset which comprises various medical feature modalities, to predict a diverse and large range of clinical tasks beyond traditional predictive scenarios, such as cardiac and non-cardiac diagnoses, as well as patient deterioration such as mortalities, ICU admission, organ-specific failure, and mechanical ventilation to name a few.
@@ 2.1 Clinical workflow and dataset creation
Figure 1 presents a schematic illustration of the proposed MDS-ED pipeline which task focuses on a specific workflow with high clinical relevance, where for each ED visit, it encompasses a set of features collected from a window of 90 minutes from the patient’s arrival at the ED to predict patient discharge diagnoses and deterioration through the stay. MDS-ED was created by linking ECG waveforms from the MIMIC-IV-ECG [18] dataset to clinical features and outcomes as clinical ground truth from the MIMIC-IV and MIMIC-IV-ED dataset [19].
@@
In this work, we broadly categorize all prediction tasks into two groups: patient discharge diagnoses and patient deterioration. For discharge diagnoses, we follow MIMIC-IV-ECG-ICD [11], which proposed a framework to predict discharge diagnoses based on a single electrocardiogram (ECG). To this end, we frame the task as a multilabel classification task in terms of ICD-10 CM codes. As in [11], we convert ICD-9 to ICD-10 CM codes where necessary, convert codes to 5 digits and propagate codes to parent codes up to the third digit. In this way, we obtain a total set of 1428 significantly populated ICD-10 CM codes covering a wide range of cardiac and non-cardiac conditions. For patient deterioration, we build on a consensus definition [20] and investigate 15 deterioration events in a multilabel setting after the patient arrived at different intervals and after the 1.5-hour feature collection window: Six subtasks cover clinical deterioration within 24 hours and cover severe hypoxemia, extracorporeal membrane oxygenation (ECMO), use of vasopressors, use of inotropes, mechanical ventilation, and in-hospital cardiac arrest (IHCA). Two subtasks cover ICU admission within 24 hours and or within the entire stay, and finally, 7 subtasks cover mortality prediction at different horizons ranging from inpatient mortality, 24 hours, 7 days, 28 days, 90 days, 180 days, to 365 days. With these two main groups, we aim to cover a wide range spectrum of patient monitoring in the ED as well as long-term mortality across 1,443 unique target labels, see Table 5 in the supplementary material for a detailed target definition discussion.
@@
While aiming to include mainly data at triage, we also include irregularly sampled vital parameters and lab values captured during the 1.5h data collection window. To capture trends, we compute simple sample-wise statistical aggregation functions per vital parameter and laboratory values, namely mean, median, minimum, maximum, standard deviation, first and last values, rate of change of them between first and last value, as well as the slope of a linear model fitted on the minutes difference between value collection and arrival as independent variable and the actual values as dependent variables. We follow [21] to select an appropriate set of laboratory values that avoids a bias towards particular prediction targets. We obtain a total of 470 features across all feature modalities in a tabular format plus a 12-lead 10-second ECG waveform per sample.
MDS-ED consists of data from 71,098 patients, encompassing 121,195 unique visits, and a total of 129,095 samples. We leverage the stratified splits provided along with the dataset accompanying [11], which includes stratification based on gender, age bins, and discharge diagnoses. We distribute the total number of 20 stratified folds into train, validation, and test following ratios of 18:1:1, see Appendix A in the supplementary material for further preprocessing and definitions details as well as descriptive statistics on demographics, biometrics, vital parameters and laboratory values in Tables 6-8.

### v3

# a Benchmark Dataset for Diagnoses and Deterioration Prediction in Emergency Medicine

@@ Abstract
Background: A clinically meaningful comparative assessment of medical decision support in emergency care is challenging due to a lack of appropriate datasets with multimodal input modalities and comprehensive prediction task. This hampers measurable progress in the field.
Results: We introduce a dataset based on MIMIC-IV, a benchmarking protocol, and initial results for evaluating multimodal decision support in the emergency department (ED). We use diverse data modalities from the first 1.5 hours after patient arrival, including demographics, biometrics, vital signs, lab values, and electrocardiogram waveforms. We analyze 1443 clinical labels across two contexts: predicting diagnoses and patient deterioration. Our diagnostic model achieves an AUROC score over 0.8 in a statistically significant manner for 609 out of 1428 conditions, including cardiac conditions like myocardial infarction and non-cardiac conditions such as renal disease and diabetes. The deterioration model scores above 0.8 in a statistically significant manner for 14 out of 15 targets, including critical events like cardiac arrest, mechanical ventilation, intensive care unit admission, as well as short- and long-term mortality. Furthermore, we provide one of the first robust demonstrations of the significant impact of raw waveform input data on model performance.
Conclusions: This study highlights the proposed dataset as a unique resource to foster progress towards measurable progress in the domain of algorithmic decision support in emergency care. The presented multimodal baseline models showcase the potential of diagnostic decision support in the field and provide strong incentives for including raw waveform data.
keywords:
Medical Decision Support; Multimodal Data; Emergency Department; Diagnostic Prediction; Patient Deterioration.
journal: general
{keypoints*}
  • Introduction of a novel dataset based on MIMIC-IV for benchmarking multimodal decision support in the emergency department.
  • Evaluation of 1443 clinical labels for diagnostic and patient deterioration predictions, achieving high AUROC scores.
  • Demonstration of significant performance improvements by incorporating raw ECG waveform data.
  • Highlighting the dataset’s comprehensive feature set collected within the first 1.5 hours after patient arrival.
  • Potential impact on improving decision-making processes in acute and emergency medicine.
@@ 1 Introduction
Artificial intelligence (AI) is an emerging field that has significantly enhanced healthcare and medicine [1] in areas such as precision medicine [2] and drug discovery [3]. The implementation of AI models is particularly relevant in acute and emergency medicine, where clinicians must address critical conditions within short time frames to make optimal clinical decisions. AI models can improve acute care by enabling early diagnosis [4], proposing tailored diagnostic workups [5], predicting admissions (including ward or intensive care unit (ICU)) [6], estimating survival rates [7], and providing faster and low-cost diagnoses [8, 9]. There has been exponential growth in the number of scientific publications on new AI models. However, many of these studies have limitations, such as predicting only a limited set of specific conditions [10, 4], offering short time horizons for survival predictions, and requiring costly diagnostic tests [11].
@@
Healthcare datasets are essential for advancing medical research and innovation. However, much of this valuable data remains undisclosed due to commercial interests and privacy concerns. Despite these hurdles, there has been a recent surge in publicly accessible datasets aimed at facilitating the development and validation of machine learning models to tackle complex healthcare challenges [12, 13, 14, 15, 16]. Nonetheless, these datasets still have limitations, including size constraints [13], lack of open-source availability [17, 18], and narrow scopes in terms of prediction tasks [13, 14, 18, 15, 17, 16]. Benchmarking is crucial for evaluating the performance and reliability of novel AI-models, ensuring their applicability and robustness in real-world clinical settings. In this study, we investigate multimodal decision support for emergency department (MDS-ED) using an open-source biomedical multimodal dataset which comprises various medical feature modalities collected within the first 1.5 hours of patient arrival at the ED department such as demographics, biometrics, vital parameters trends, laboratory trends, and electrocardiogram (ECG) waveforms, to predict a diverse and large range of clinical tasks beyond traditional predictive scenarios, such as cardiac and non-cardiac diagnoses, as well as patient deterioration such as mortalities, ICU admission, organ-specific failure, and mechanical ventilation to name a few.
2 Analyses
2.1 Benchmarking predictive performance
Table 1: Predictive performance (macro AUROC with 95% bootstrap confidence intervals) across all scenarios for unimodal models as well as multimodal prediction models. Where the clinical routine data encompasses a comprehensive set of 470 variables from diverse feature modalities such as demographics, biometrics, vital parameters trends, and laboratory values trends. All the scenarios can also be split into model architecture, either tree-based or deep-learning models.
Diagnoses Deterioration
Unimodal models
Clinical routine data (tree-based) 0.7473 (0.7459, 0.7484) 0.8365 (0.8179, 0.8546)
ECG features only (tree-based) 0.6366 (0.6352, 0.6374) 0.7385 (0.7114, 0.7588)
ECG waveforms only (deep-learning) 0.7678 (0.7636, 0.7718) 0.8565 (0.8410 , 0.8707)
Multimodal models
ECG features + clinical routine data (tree-based) 0.7685 (0.7676, 0.7705) 0.8656 (0.8498, 0.8830)
ECG waveforms + clinical routine data (deep-learning) 0.8256 (0.8222,0.8288) 0.9115 (0.8991,0.9222)
Table 1 presents the predictive performance of both unimodal and multimodal models across the diagnoses and deterioration settings. For diagnoses prediction, the multimodal model that integrates ECG waveforms and clinical routine data achieved a macro average across all areas under the respective receiver operating curves (AUROC) (macro-AUROC) of 0.8256 (0.8222,0.8288), which is notably higher than the rest of the models. Similarly, in the deterioration prediction task, the multimodal model that integrates ECG waveforms and clinical routine data achieved an AUROC of 0.9115 (0.8991,0.9222), which also overcomes the rest of the models. In the following discussion, we restrict ourselves to the discussion of the most comprehensive multimodal model based on ECG waveforms and clinical routine data.
2.2 Task-dependent predictive performance
Table 2: Best-performing individual statements organized according to selected ICD chapters underscoring the breadth of accurately predictable statements. The table shows the four best-performing individual statements per ICD chapter (8 for chapter IX (Circulatory system diseases), 4 for the rest), where we show only AUROC scores above 0.85 and where also the lower bound of the 95% bootstrap confidence interval exceeds 0.80. To showcase the breadth of reliably predictable statements, we list only the best-performing statement per 3-digit ICD code.
IX: Circulatory system. 0.8761 (126/181 >0.80 lower bound)
I071: 0.9461 [287] Rheumatic tricuspid insufficiency I080: 0.8522 [770] Rheumatic mitral and aortic valve disorders
I110: 0.9044 [3185] Hypertensive heart disease with heart failure I120: 0.9756 [3349] Hypert. CKD (stage 5 or with end-stage renal disease)
I130: 0.9426 [2994] Hypertensive heart and kidney disease with heart failure and CKD I160: 0.9524 [427] Hypertensive urgency
I200: 0.8934 [810] Unstable angina I2109: 0.9734 [401] STEMI anterior wall coronary artery
II: Neoplasms. 0.8572 (34/53 >0.80)
C228: 0.9628 [235] Malignant neoplasm of liver, primary C259: 0.8991 [275] Malignant neoplasm of pancreas
C341: 0.8739 [463] Malignant neoplasm of upper lobe or lung C5091: 0.8544 [425] Malignant neoplasm of breast, female
XIX: Injury, poisoning, and certain other consequences of external causes. 0.8479 (37/86 >0.80)
S0100: 0.9161 [258] Unspecified open wound of scalp S065: 0.8767 [783] Traumatic subdural hemorrhage
S12: 0.9356 [392] Fracture of cervical vertebra and other parts of neck S2249: 0.9198 [409] Multiple fractures of ribs
I: Certain infectious and parasitic diseases. 0.8346 (19/52 >0.80)
A0472: 0.9374 [235] Enterocolitis due to Clostridium difficile A084: 0.8709 [352] Viral intestinal infection,
A40: 0.9347 [339] Streptococcal sepsis A4151: 0.9522 [626] Sepsis due to Escherichia coli [E. coli]
IV: Endocrine, nutritional, and metabolic diseases. 0.8336 (52/107 >0.80)
E1022: 0.9661 [309] Type 1 diabetes mellitus with diabetic CKD E1121: 0.9466 [707] Type 2 diabetes mellitus with diabetic nephropathy
E1342: 0.9271 [1809] Diabetes mellitus with diabetic polyneuropathy E43: 0.9329 [1417] Severe protein-calorie malnutrition
III: Diseases of the blood and blood-forming organs. 0.8335 (20/45 >0.80)
D500: 0.8557 [897] Cronic Iron deficiency anemia D6181: 0.8761 [1503] Pancytopenia
D62: 0.8843 [4514] Acute posthemorrhagic anemia D630: 0.8920 [589] Anemia in neoplastic disease
XI: Diseases of the digestive system. 0.827 (40/119 >0.80)
K264: 0.9189 [255] Chronic duodenal ulcer with hemorrhage K3181: 0.9384 [419] Angiodysplasia of stomach and duodenum
K521: 0.9290 [257] Toxic gastroenteritis and colitis K566: 0.8958 [598] Other and unspecified intestinal obstruction
XIV: Diseases of the genitourinary system. 0.8253 (18/39 >0.80)
N08: 0.9393 [798] Glomerular disorders N12: 0.8941 [325] Tubulo-interstitial nephritis, not acute or chronic
N170: 0.9092 [1773] Acute kidney failure with tubular necrosis N183: 0.8625 [4734] Chronic kidney disease, stage 3
X: Diseases of the respiratory system. 0.8226 (33/71 >0.80)
J1521: 0.9202 [275] Pneumonia due to staphylococcus aureus J39: 0.9144 [276] Other diseases of upper respiratory tract
J440: 0.8936 [440] Chronic obstructive pulmonary disease with acute lower respiratory infection J690: 0.8968 [2231] Pneumonitis due to inhalation of food and vomit
V: Mental, behavioral and neurodevelopmental disorders. 0.8111 (32/71 >0.80)
F0280: 0.8556 [1540] Dementia F0390: 0.8565 [3690] Unspecified dementia
F1012: 0.9591 [625] Alcohol abuse with intoxication F1110: 0.8735 [666] Opioid abuse, uncomplicated
XII: Diseases of the skin and subcutaneous tissue. 0.8087 (10/25 >0.80)
L0311: 0.8646 [573] Cellulitis of other parts of limb L408: 0.8933 [292] Other psoriasis
L8915: 0.8920 [1209] Pressure ulcer of sacral region L974: 0.9093 [336] Non-pressure chronic ulcer of heel and midfoot
VI: Diseases of the nervous system. 0.7955 (19/67 >0.80)
G309: 0.8508 [1192] Alzheimer’s disease G4030: 0.9374 [244] Idiopathic epilepsy and epileptic syndromes
G609: 0.8772 [816] Hereditary and idiopathic neuropathy G8190: 0.9251 [265] Hemiplegia
XVIII: Symptoms, signs and abnormal clinical and laboratory findings 0.7813 (35/171 >0.80)
R000: 0.8507 [2714] Tachycardia, unspecified R042: 0.8745 [316] Hemoptysis
R0600: 0.8502 [1999] Dyspnea R1032: 0.8956 [453] Left lower quadrant pain
XIII: Diseases of the musculoskeletal system and connective tissue. 0.7405 (14/68 >0.80)
M100: 0.9106 [246] Idiopathic gout M129: 0.8933 [495] Arthropathy
M1A9: 0.9000 [253] Chronic gout, unspecified M810: 0.8439 [4398] Age-related osteoporosis without fracture
Table 2 demonstrates the diagnoses that can be predicted most accurately by the model, which we organized based on ICD chapters. The algorithm’s predictive performance ranges from an AUROC of 0.7405 for the musculoskeletal system and connective tissue (XIII) chapter to 0.8761 for the circulatory system (IX) chapter. In total, the model predicted 609 out of 1,428 individual ICD diagnoses with high accuracy, which we define in this work as conditions where the lower bound of the 95% bootstrap confidence interval for AUROC exceeds 0.80.
Label: AUROC [instances] Label: AUROC [instances]
Clinical deterioration. 0.9070 (5/6 > 0.80) lower bound
Severe hypoxemia: 0.6980 [555] ECMO: 0.9355 [166]
Vasopressors: 0.9239 [1171] Inotropes: 0.9400 [375]
Mechanical ventilation: 0.9590 [4223] IHCA: 0.9859 [623]
ICU admission. 0.9063 (2/2 > 0.80)
ICU 24 hours: 0.9147 [15677] ICU overall: 0.8979 [18860]
Mortality. 0.9168 (7/7 > 0.80)
In-hospital: 0.9423 [99] 24 hours: 0.9600 [816]
7 days: 0.9429 [2273] 28 days: 0.9115 [5150]
90 days: 0.8952 [9297] 180 days: 0.8894 [12839]
365 days: 0.8768 [17756]
Table 3 showcases the deterioration model’s predictive performance across clinical deterioration, ICU admission, as well as both short and long-term mortality with many AUROC scores significantly exceeding the 0.80 threshold. Overall, for clinical deterioration, the model achieves an AUROC of 0.9070, whereas, for ICU admissions, the model reports an overall AUROC of 0.9063, and for mortality predictions also exhibits high performance, with an overall AUROC of 0.9168.
3 Discussion
3.1 Impact of data modalities
From the results presented in the previous section, we can draw several conclusions: Firstly, the results demonstrate that multimodal models, which integrate diverse data types, offer superior performance in both diagnostic and deterioration tasks (row 4&5 vs. the rest). Secondly, in the diagnoses task as well as in the deterioration task, the use of ECG raw waveforms instead of ECG features improves the performance in a statistically significant manner (row 4 vs. row 5), finding which is not in line with [14]. To the best of our knowledge, this is the first statistically robust demonstration of the added value of raw ECG waveform input against ECG features for clinically relevant prediction tasks such as diagnoses and deterioration prediction. Thirdly, for the unimodal models, in the deterioration task, the clinical routine data model outperforms ECG features only and ECG waveforms only, however, in the diagnoses task, the ECG waveforms only outperforms the other 2 settings, we hypothesize that for the deterioration task, the clinical routine data apart of including a rich set of clinical features (demographics, biometrics, vital parameters trends, and laboratory values trends) against only an single ECG either in features or waveform, it also includes trends over time which aligns with the task definition of deterioration. Despite this, we believe that a single ECG snapshot can achieve high performances for both tasks, but also we believe that the inclusion of multiple ECGs over time instead of just a single snapshot would allow us to capture more meaningful deterioration and potentially diagnoses trends.
3.2 Clinical significance
From a clinical point of view, we conclude that our diagnostic models exhibit high predictive capabilities for acute, chronic, as well as management conditions. Acute conditions like ST elevation myocardial infarction, acute pericarditis, persistent atrial fibrillation, and acute kidney failure require rapid diagnosis and immediate intervention to prevent severe outcomes. The model’s strong performance in identifying these conditions highlights its potential to significantly impact patient care through timely and accurate diagnosis. Conversely, chronic conditions such as hypertensive heart and chronic kidney disease, ischemic cardiomyopathy, dilated cardiomyopathy, and end-stage renal disease benefit from ongoing management and long-term treatment strategies. The model’s ability to effectively predict these chronic conditions underscores its utility in supporting sustained patient monitoring and management, ultimately improving patient outcomes through consistent and reliable diagnostic support. Finally, detecting management conditions like thrombosis of prosthetic devices enhances ongoing care and ensures timely intervention. As demonstrated in [12], even the model based only on ECG waveforms can reliably predict cardiac but most notably non-cardiac conditions. This applies even more to the multimodal modal that involves clinical features as well as ECG waveforms.
Our deterioration model has major clinical relevance by enhancing early recognition of patient deterioration and supporting physician decisions in the emergency department. Its high predictive performance across various scenarios effectively identifies patients at risk for severe complications, which is crucial in acute and emergency medicine. The model’s ability to reliably predict clinically relevant indicators, such as vasopressor use, inotrope use, mechanical ventilation, and in-hospital cardiac arrest, enables healthcare providers to anticipate these events and implement early, tailored interventions. Additionally, its robust performance in predicting ICU admissions allows for better resource allocation, ensuring that patients who need intensive care receive it promptly while preventing unnecessary admissions. The model accurately predicts 24-hour in-hospital mortality and accurately predicts 365-day mortality, supporting clinical decision-making in prioritizing high-risk patients and discussing prognosis with families. Long-term mortality predictions aid in advanced care planning, such as deciding against ICU admission for patients with a short expected lifespan after discharge, and inform post-discharge planning and follow-up care to prevent readmissions and improve long-term health outcomes.
3.3 Limitations
Despite the promising results, there are several limitations to consider. Firstly, the presented models rely on the specifics of Beth Israel Deaconess Medical Center in terms of patient demographics, available resources, process characteristics, and data collection practices. We carefully quantify the statistical uncertainties of our results and are convinced of the robustness of our results for this particular scenario within the limits of a retrospective evaluation. However, we cannot make any statements about the transferability of our findings to other hospitals or deployment scenarios as this would require an external validation dataset, see [19] for a recent perspective on the role of external validation for clinical decision support. Secondly, the reliance on data collected within the first 1.5 hours of the ED visits may limit the model’s applicability. We deliberately decided to focus on this setting to keep the dataset homogeneous. However, there is also a clear clinical case for a prediction model only incorporating information available during triage. A similar dataset to the one proposed could be created, but this task is left for future research. Thirdly, while we focus on the construction of a benchmark dataset and provide the first results for it, we do not provide any insights into the trained models beyond a discussion of the model performance for specific conditions. This could be achieved for example through attribution methods proposed in the context of explainable AI, however, mostly in uni-modular settings [20], to verify that the models do not exploit spurious correlations to reach their strong performance. Again, a detailed investigation also of causal vs. associational attributions along the lines of [21] is beyond the scope of this submission. Finally, although the open-source dataset fosters innovation, the implementation of a large-scale single end-to-end model with a wide range of target variables such as the diagnoses case might result in challenges during deployment in resource-constrained hardware, see 9 in the supplementary material for a more extensive discussion.
3.4 Future research directions
4 Potential implications
@@ 5.1 Clinical workflow and dataset creation
Figure 1 presents a schematic illustration of the proposed MDS-ED pipeline which task focuses on a specific workflow with high clinical relevance, where for each ED visit, it encompasses a set of features collected from a window of 90 minutes from the patient’s arrival at the ED to predict patient discharge diagnoses and deterioration through the stay. MDS-ED was created by linking ECG waveforms from the medical information mart for intensive care (MIMIC) database, specifically, MIMIC-IV-ECG [23] dataset to clinical features and outcomes as clinical ground truth from the MIMIC-IV and MIMIC-IV-ED dataset [24].
@@
In this work, we broadly categorize all prediction tasks into two groups: patient discharge diagnoses and patient deterioration. For discharge diagnoses, we follow MIMIC-IV-ECG-ICD [12], which proposed a framework to predict discharge diagnoses based on a single electrocardiogram (ECG). To this end, we frame the task as a multilabel classification task in terms of international classification of diseases clinical modification 10 (ICD-10 CM) codes. As in [12], we convert ICD-9 to ICD-10 CM codes where necessary, convert codes to 5 digits, and propagate codes to parent codes up to the third digit. In this way, we obtain a total set of 1428 significantly populated ICD-10 CM codes covering a wide range of cardiac and non-cardiac conditions. For patient deterioration, we build on a consensus definition [25] and investigate 15 deterioration events in a multilabel setting after the patient arrived at different intervals and after the 1.5-hour feature collection window: Six subtasks cover clinical deterioration within 24 hours and cover severe hypoxemia, extracorporeal membrane oxygenation (ECMO), use of vasopressors, use of inotropes, mechanical ventilation, and in-hospital cardiac arrest (IHCA). Two subtasks cover ICU admission within 24 hours and or within the entire stay, and finally, 7 subtasks cover mortality prediction at different horizons ranging from inpatient mortality, 24 hours, 7 days, 28 days, 90 days, 180 days, to 365 days. With these two main groups, we aim to cover a wide range spectrum of patient monitoring in the ED as well as long-term mortality across 1,443 unique target labels, see Table 5 in the supplementary material for a detailed target definition discussion.
@@
While aiming to include mainly data at triage, we also include irregularly sampled vital parameters and lab values captured during the 1.5h data collection window. To capture trends, we compute simple sample-wise statistical aggregation functions per vital parameter and laboratory values, namely mean, median, minimum, maximum, standard deviation, first and last values, rate of change of them between first and last value, as well as the slope of a linear model fitted on the minutes difference between value collection and arrival as independent variable and the actual values as dependent variables. We follow [26] to select an appropriate set of laboratory values that avoids a bias towards particular prediction targets. We obtain a total of 470 features of clinical routine data across all feature modalities in a tabular format plus a 12-lead 10-second ECG waveform per sample. Similarly, the ECG feature set is composed of RR interval, P-onset, QRS-onset, QRS-end, and T-end in milliseconds (ms), whereas P-axis, QRS-axis, and T-axis in degrees.
MDS-ED consists of data from 71,098 patients, encompassing 121,195 unique visits, and a total of 129,057 samples. We leverage the stratified splits provided along with the dataset accompanying [12], which includes stratification based on gender, age bins, and discharge diagnoses. We distribute the total number of 20 stratified folds into train, validation, and test following ratios of 18:1:1. For deep learning models with clinical routine data, we apply a median imputation learned from train set to validation and test for each column and we apply a masking strategy in which we add an additional binary column to indicate which values were imputed to the model. See the supplementary material for further preprocessing and definitions details as well as descriptive statistics on demographics, biometrics, vital parameters and laboratory values in Tables 6-8.